Circular No. 6605

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bl IN ANONYMOUS GALAXY Delphine Hardin reports the discovery on Mar. 7 of a supernova (mag

V ∼ 21.5), found in the course of a systematic CCD search by the EROS team with the Marly 1-m telescope at the European Southern Observatory. SN 1997bl is located at α = 12^{h}40^{m}21^{s}.0, δ = −11^{o}44′39′′ (equinox 2000.0), which is 2′′ from the center of the host galaxy. A spectrogram obtained by M. T. Ruiz on Mar. 16 with the ESO 3.6-m telescope indicates this to be a type-I supernova.

GS 1843+00 D. Chakrabarty, Massachusetts Institute of Technology; T. Takeshima,

Goddard Space Flight Center, NASA, and Universities Space Research Association; and R. B. Wilson, Marshall Space Flight Center, NASA, report: “Scans of the region containing this accreting x-ray pulsar (IAUC 6585, 6595), using the PCA detectors on the RXTE spacecraft on Mar. 16.04 UT, have yielded the following improved position for the x-ray source: α = 18^{h}45^{m}.6, δ = +0^{o}50′ (equinox J2000.0). The error radius of 2′ (90-percent confidence) is dominated by systematic uncertainty due to source-intensity variations during the scans. Optical observations of this region to identify the mass donor are encouraged. Observations with the RXTE All Sky Monitor and the BATSE detectors on the Compton Gamma Ray Observatory indicate that the x-ray source was still active as of Mar. 23.”

GRO J0422+32 A. Iyudin and F. Haberl, Max-Planck-Institut für Extraterrestrische

Physik, report: “The x-ray nova known as GRO J0422+32 (IAUC 5580, 5582) was detected again by the All-Sky Monitor (ASM) on RXTE during the time period 1996 Nov. 18–1997 Jan. 15. The source flux was 4.4 mCrab in the energy band 2–10 keV, assuming a Crab-like power-law spectrum. The data were analyzed from the publicly available ASM quick-look results

provided by the ASM/RXTE team (http://heasarc.gsfc.nasa.gov/

docs/xte/ASM). The repeated outburst of GRO J0422+32 happened at a time consistent with the expected outburst period of ∼ 120 days, suspected on the basis of earlier x-ray (BATSE, ASCA; cf. IAUC 5851) and optical measurements (Chevalier and Ilovaisky 1995, A.Ap. 297, 103). The repetition of the nova outburst indicates similarities between GRO J0422+32 and low-mass binary systems such as Aql X-1.”

1997 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green